If the probability of a student taking a math class is 0.70, the probability of taking a physics class is 0.30, and the probability of taking a math class and a physics class is 0.15, what is the probability of a student taking a math class or a physics class?
Answer by ikleyn(53763)

.

    p(M or P) = p(M) + p(P) - p(M and P) = 0.7 + 0.3 - 0.15 = 0.85.    ANSWER

Solved, answered and explained.